# Knowledge Integration Strategies in Autonomous Vehicle Prediction and Planning: A Comprehensive Survey

### v1

@@ Abstract
This comprehensive survey examines the integration of knowledge-based approaches into autonomous driving systems, with a focus on trajectory prediction and planning. We systematically review methodologies for incorporating domain knowledge, traffic rules, and commonsense reasoning into these systems, spanning purely symbolic representations to hybrid neuro-symbolic architectures. In particular, we analyze recent advancements in formal logic and differential logic programming, reinforcement learning frameworks, and emerging techniques that leverage large foundation models and diffusion models for knowledge representation. Organized under a unified literature survey section, our discussion synthesizes the state-of-the-art into a high-level overview, supported by a detailed comparative table that maps key works to their respective methodological categories. This survey not only highlights current trends—including the growing emphasis on interpretable AI, formal verification in safety-critical systems, and the increased use of generative models in prediction and planning—but also outlines the challenges and opportunities for developing robust, knowledge-enhanced autonomous driving systems.
Index Terms:
Formal Logic, Knowledge Representation, Knowledge Integration, Conflict Resolution, Literature Survey, Planning, Prediction, LLM, Survey
@@ I Introduction
Autonomous driving has evolved into a complex and multifaceted research field. Vehicles’ safe and efficient operation relies on accurately anticipating and planning motions in real-time. While significant progress has been made in simulation environments, the true challenge lies in transitioning automated vehicles from controlled simulations to the unpredictable real world. In real-world scenarios, vehicles must contend with a wide variety of situations, including unforeseen behaviors and novel circumstances. While effective in structured environments, rule-based systems often require adaptation to accommodate dynamic traffic conditions. On the other hand, machine learning (ML)-based solutions, though powerful, typically demand extensive amounts of data and can struggle when deployed in out-of-distribution scenarios or unfamiliar domains. Unlike machines, human drivers excel in such situations by leveraging a combination of rules, knowledge, and general driving intuition to make informed decisions. Integrating such human-like reasoning—infusing rules and domain knowledge—into the autonomous driving pipeline could enhance the system’s accuracy, robustness, and ability to handle new or unexpected scenarios. Two critical components in this domain are trajectory prediction and trajectory planning. In our context, trajectory prediction refers to the process of forecasting the future paths of other traffic agents based on observed historical data and contextual information such as road geometry, dynamic interactions, and regulatory constraints. Trajectory planning, on the other hand, involves generating a feasible and optimal motion path for the controlled (ego) vehicle that satisfies safety, legal, and comfort constraints while taking into account the anticipated behavior of other road users.
Despite numerous surveys addressing various aspects of autonomous driving—from computer vision and object detection to deep reinforcement learning and semantic segmentation—there remains a notable gap in comprehensive reviews focusing specifically on integrating explicit knowledge representations into trajectory prediction and planning. In particular, there is a lack of surveys examining recent foundation and diffusion model–based methodologies. These emerging approaches have gained traction due to their ability to leverage vast amounts of unstructured data and advanced generative techniques to produce precise and adaptable predictions. They can embed traffic rules and environmental constraints either implicitly within learned representations or explicitly through integrated symbolic reasoning, thereby enhancing both interpretability and performance in complex driving scenarios.
This survey is driven by several key observations:
@@
  • Fragmented Literature: Previous reviews have explored trajectory prediction and planning [1]. However, they rarely holistically integrate these tasks under frameworks combining traditional methods with modern generative techniques and extensive domain knowledge. We examine how knowledge is integrated into the automated driving domain in various ways, providing a unified perspective.
@@
  • Emerging Methodologies: Recent advances, including foundation model-based approaches (e.g., large language models and retrieval-augmented generation) and diffusion model-based methods, offer promising new avenues for integrating explicit regulatory and spatial constraints into prediction and planning. These methods can revolutionize how autonomous systems interpret contextual information and generate compliant, safe trajectories.
@@
  • Practical and Future Relevance: In real-world autonomous driving, it is crucial that prediction and planning modules not only operate accurately but also adhere to traffic rules and adapt dynamically to complex environments. The integration of explicit knowledge—whether embedded implicitly through data-driven models or explicitly via hybrid symbolic frameworks—can significantly enhance system robustness and safety.
@@
We focus exclusively on works that explicitly integrate rules or knowledge. For instance, [2] utilized road boundary information in a CNN model but did not explicitly incorporate traffic rules, so we exclude such works. By offering a detailed taxonomy and critical analysis of state-of-the-art methods—including classical approaches and the latest foundation and diffusion model–based techniques—this survey aims to provide researchers and practitioners with a unified perspective on achieving precise, rule-compliant trajectory prediction and planning.
By critically evaluating both symbolic and hybrid methodologies, we seek to offer insights that can drive the development of next-generation autonomous systems—systems that are not only capable of understanding and predicting dynamic traffic scenarios but are also inherently safe, interpretable, and aligned with regulatory requirements. Autonomous driving relies on robust understanding and reasoning about driving environments, traffic rules, and commonsense knowledge. This survey explores knowledge representation methods for prediction and planning in automated driving, covering knowledge-graph and ontology-based, hybrid, reinforcement learning, LLM-based, and formal logic-based approaches. We aim to provide a comprehensive overview for researchers and practitioners, fostering advancements in knowledge-driven autonomous driving.
@@
Reinforcement Learning (RL) [17, 18, 19, 20, 21, 22, 23] Embeds traffic rules in state representations and reward functions. Employs hybrid strategies (e.g., imitation learning, formal logic) to improve safety and exploration. Advantages: Adaptive policy learning from dynamic environments. Limitations: Reward engineering complexity and sample inefficiency.
LLM & RAG Methods [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] Uses unstructured text and regulatory documents to adapt driving behavior. Translates natural language rules into formal (temporal) logic and generates context-aware decisions. Often combined with multimodal inputs (e.g., DriveGPT4). Advantages: Dynamic adaptation and enhanced interpretability. Limitations: Latency concerns and risk of LLM hallucinations.
Formal Logic-Based Methods [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] Formalizes traffic regulations via temporal logic, ASP, etc. Integrated into verification, MPC, and optimization frameworks for provable safety. Advantages: Strong safety guarantees and high interpretability. Limitations: High computational complexity and challenges with real-time integration.
Hybrid Symbolic & Neural Methods [36, 40] (also in Formal Logic), [49, 50, 51, 52, 53, 54, 55, 56, 57, 58] [32] (also in LLM) Fuses deep learning–based perception with explicit symbolic reasoning. Incorporates symbolic constraints in loss functions or via LLM–generated rules. Advantages: Balances flexibility with rule compliance. Limitations: Increased integration complexity and synchronization challenges.
Diffusion & Other Learning–Based Approaches [59, 60] Leverages diffusion processes to generate multimodal, rule–compliant trajectories. Blends stochastic exploration with explicit constraint encoding. Advantages: Captures diverse trajectory distributions. Limitations: Balancing exploration with strict rule adherence can be challenging.
@@ II Literature Survey
This section presents a high-level overview of the diverse methodologies proposed for integrating domain knowledge, traffic rules, and commonsense reasoning into autonomous driving systems. Our survey is organized into several subsections—ranging from Knowledge-Graph and Ontology-Based Methods, Reinforcement Learning (RL) methods, and Large Language Model (LLM) approaches to Formal Logic-Based and Hybrid Neuro-Symbolic Methods—that collectively cover the spectrum of current research. To guide the reader through these varied approaches, Table I provides a compact, comparative summary of the key works discussed in this survey. The table maps each representative study to its corresponding methodological category, highlights its core contributions, and outlines its advantages and limitations. Also, note that some works fall into multiple categories due to their interdisciplinary nature. This high-level synthesis facilitates an understanding of how these diverse approaches interconnect and serves as a roadmap for the detailed discussions that follow in the subsequent sections.
@@ II-A Knowledge-Graph and Ontology-Based Methods
Recent advancements in autonomous driving have highlighted the importance of structured knowledge representation using ontologies and knowledge graphs (KGs) to model complex urban environments, enhance prediction accuracy, and enable rule-compliant planning. These frameworks provide explicit semantic representations of static road elements (e.g., lanes, traffic signs), dynamic interactions (e.g., vehicle-pedestrian relationships), and regulatory constraints, enabling interpretable reasoning across heterogeneous data sources.
@@
Early works demonstrated the potential of ontology-based reasoning for driving scenarios. [3] pioneered a scenario-adaptive approach using Web Ontology Language (OWL) to formalize spatial states and vehicle-road relationships, integrating these with hidden Markov models for maneuver prediction. At the same time, their method improves prediction horizons through domain knowledge and manual ontology engineering limits scalability. Similarly, [4] developed intersection-specific ontologies with rule-based decision-making, and [5] established a framework for querying static and dynamic traffic semantics in real-time. These studies underscore the trade-off between interpretability and adaptability: rule-based ontologies excel in structured environments like intersections [4] but struggle with unseen scenarios.
@@
SemanticFormer [6] improves autonomous driving trajectory prediction by considering traffic participants, road topology, and signs, enhancing performance by 4-5% when added to existing trajectory predictors. Hybrid methods combining ontologies with probabilistic models address uncertainty inherent in traffic interactions. [7] merged high-level traffic coordination ontologies with low-level trajectory models, while [8] integrated description logic with probabilistic semantics to handle incomplete information. [9] extended this by encoding interaction probabilities for long-term behavior prediction, though computational delays hindered real-time control. [10] augmented scene understanding by integrating KGs extracted from textual corpora, enabling causal reasoning for novel scenarios. However, offline knowledge curation in their approach limits responsiveness to dynamic environments.
@@
Ontologies also support real-time decision-making through hierarchical reasoning. [11] leveraged driving experience knowledge bases for situation estimation, while [12] modeled critical factors via joint description logic and rule-based reasoners. These methods enhance transparency in regulatory compliance, but face scalability challenges as ontology complexity grow [5]. A comprehensive survey by [13] emphasizes that while ontological frameworks improve interoperability and rule awareness, their rigidity complicates adaptation to unexpected events.
@@
Current research focuses on mitigating manual engineering burdens through standardization and automation. Domain-specific ontologies like ASAM OpenXOntology [^1] https://www.asam.net/standards/asam-openxontology/ provide reusable templates for traffic concepts, reducing development efforts [14, 15]. Promisingly, Large Language Models (LLMs) are being explored for semi-automated KG generation [16], potentially enabling dynamic knowledge integration with sensor data. Such advances aim to balance the interpretability of symbolic reasoning with the adaptability of data-driven methods, addressing long-standing challenges in real-time performance and scalability [13, 9].
@@ II-B Reinforcement Learning (RL) Methods
Reinforcement learning (RL) has emerged as a powerful paradigm for developing autonomous driving systems, enabling sequential decision-making through dynamic environment interactions. By integrating safety constraints, regulatory compliance, and structured domain knowledge into reward functions and state representations, RL frameworks excel in learning adaptive policies for complex driving scenarios. Recent advancements in this domain focus on curriculum learning, safety-centric reward design, and sim-to-real transfer while addressing challenges such as sample inefficiency and reward engineering.
A key trend in RL-based methods is using structured state encodings to embed traffic rules and safety priors. For instance, [17] propose a graph-based RL framework that employs curriculum learning to increase the complexity of highway driving scenarios progressively. By integrating explicit traffic-rule information into the state representation using graph neural networks, their method accelerates learning and improves collision avoidance and lane-change efficiency in dense traffic. Similarly, [18] address the challenge of adapting to evolving traffic regulations by designing a law-adaptive backup policy. Their approach translates natural language traffic laws into formal constraints, enabling the RL agent to recognize potential violations and trigger re-planning. While these methods enhance compliance and safety, they introduce reward engineering and policy evaluation complexity.
@@
Bridging the gap between simulation and real-world deployment remains a critical challenge in RL-based autonomous driving. [22] investigate the sim-to-real transfer of RL policies, training agents in virtual environments and adapting them to real-world conditions. Their work highlights the potential of RL for robust performance in heterogeneous scenarios but also underscores the challenges of achieving reliable generalization. Surveys such as [23] provide a broader perspective, categorizing RL-based approaches by state representation (e.g., raw sensor inputs vs. graph abstractions) and reward formulations. These surveys emphasize the trade-offs between sample efficiency and safety guarantees, offering insights into the design of scalable and reliable RL frameworks.
Despite their advantages, RL-based methods face several challenges. Reward engineering remains a non-trivial task, as balancing safety, compliance, and efficiency in reward design is critical for policy performance. Sample inefficiency is another significant bottleneck, as complex environments demand extensive training data, prolonging development cycles. The sim-to-real gap further complicates deployment, with discrepancies between simulated and real-world dynamics hindering reliable policy transfer. Additionally, the ”black-box” nature of RL policies complicates verification in safety-critical contexts, necessitating interpretable and verifiable representations.
Emerging directions in RL-based autonomous driving focus on addressing these challenges. Evolutionary strategies, such as those proposed by [19], offer efficient exploration mechanisms for safe policy learning. Formal methods like linear temporal logic (LTL) [20] provide verifiable compliance guarantees, while hybrid architectures blending imitation learning with RL [21] enhance human-like behavior and adaptability. Future research should prioritize scalable reward automation, robust sim-to-real pipelines, and interpretable policy representations to advance the real-world applicability of RL-based autonomous driving systems.

### v2

@@ Abstract
This comprehensive survey examines the integration of knowledge-based approaches in autonomous driving systems, specifically focusing on trajectory prediction and planning. We extensively analyze various methodologies for incorporating domain knowledge, traffic rules, and commonsense reasoning into autonomous driving systems. The survey categorizes and analyzes approaches based on their knowledge representation and integration methods, ranging from purely symbolic to hybrid neuro-symbolic architectures. We examine recent developments in logic programming, foundation models for knowledge representation, reinforcement learning frameworks, and other emerging technologies incorporating domain knowledge. This work systematically reviews recent approaches, identifying key challenges, opportunities, and future research directions in knowledge-enhanced autonomous driving systems. Our analysis reveals emerging trends in the field, including the increasing importance of interpretable AI, the role of formal verification in safety-critical systems, and the potential of hybrid approaches that combine traditional knowledge representation with modern machine learning techniques.
{textblock}
0.82(0.09,0.93) ©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
Accepted for publication in Proceedings of the IEEE Intelligent Vehicles Symposium (IV), Cluj-Napoca - Romania, 22-25 June 2025.
@@ I Introduction
In modular automated driving, accurately predicting the maneuvers of traffic participants and planning safe driving actions are crucial. While significant progress has been made in simulation environments, the challenge lies in transitioning automated vehicles to the unpredictable real world, where they have to handle unforeseen behaviors and novel situations. Rule-based systems perform well in structured environments but struggle with dynamic traffic conditions. Machine learning (ML)- based approaches typically require large datasets but struggles in out-of-distribution scenarios. In contrast, human drivers leverage a combination of rules, knowledge, and intuition to make informed decisions. Integrating such human-like reasoning—infusing rules and domain knowledge—into autonomous systems could enhance accuracy, robustness, and adaptability to unexpected scenarios. Two key components discussed are trajectory prediction, which uses historical data and context such as road layout, agent interactions, and traffic rules to forecast the future paths of road users. Second, trajectory planning determines an optimal and feasible path for the ego vehicle, ensuring it complies with safety, legal, and comfort standards while anticipating the predicted movements of others. Despite other surveys addressing various aspects of autonomous driving (AD), a notable gap exists in comprehensive reviews focusing specifically on integrating explicit knowledge representations into trajectory prediction and planning. Notably lacking are examinations of recent foundation and diffusion model-based methodologies, which have gained traction due to their ability to use vast amounts of unstructured data and advanced generative techniques to produce precise and adaptable predictions. In this paper, knowledge is explicit representations of rules, ontologies, priors, textual information, or symbolic constraints, and symbolic approaches are methods based on logic, rules, and ontologies without learned components. Key observations driving this survey are:
@@
  • Fragmented Literature: Previous reviews have explored trajectory prediction and planning separately [1]. Our survey uniquely examines how knowledge is holistically integrated across the automated driving domain, combining traditional methods with modern generative techniques.
@@
  • Emerging Methodologies: Recent advances in foundation models, e.g., large language models (LLMs) and retrieval-augmented generation (RAG) and diffusion models offer promising avenues for integrating explicit regulatory and spatial constraints into prediction and planning, potentially revolutionizing how autonomous systems generate compliant, safe trajectories.
@@
  • Practical Relevance: For real-world deployment, autonomous vehicles must operate accurately, adhere to traffic rules, and adapt to complex environments. Knowledge integration significantly enhances system robustness and safety, whether implicitly through data-driven models or explicitly via hybrid symbolic frameworks.
@@
We focus exclusively on works that explicitly integrate rules or knowledge, excluding approaches that utilize environmental information without incorporating traffic regulations. However, in many cases, particularly in ML-based methods, this explicit prior knowledge does not become explicit; instead, in the overall pipeline, it becomes implicit or soft constraints. This survey provides researchers and practitioners with a unified perspective on achieving precise, rule-compliant trajectory prediction and planning by offering a detailed taxonomy and critical analysis of methods spanning classical approaches to the latest foundation and diffusion model-based techniques. Our evaluation of symbolic and hybrid methodologies offers insights for developing next-generation autonomous systems that can understand and predict dynamic traffic scenarios while remaining safe, interpretable, and aligned with regulatory requirements.
I-A Related Surveys
Several recent surveys have explored various aspects of AD, but none have specifically addressed the integration of domain knowledge into both prediction and planning modules. [2] reviews architectural principles for joint prediction and planning in deep learning systems but omits explicit rule embedding. [3] analyzes end-to-end frameworks mapping raw sensor data to control commands yet excludes hybrid symbolic components. [4] surveys formal verification techniques for rule compliance, focusing primarily on offline verification rather than runtime knowledge embedding. [5] reviews knowledge graph integration in AV systems, primarily at the perception layer. [6] provides a broad overview of AV system components without a detailed taxonomy of knowledge integration approaches. In contrast, our survey delivers a unified taxonomy of knowledge integration methods spanning purely symbolic, hybrid neuro-symbolic, LLM-based, and diffusion-driven techniques, with comparative analysis of how traffic rules and domain priors are explicitly and implicitly encoded within both trajectory prediction and planning modules.
@@ I-A Related Surveys
Reinforcement Learning (RL) [21, 22, 23, 24, 25, 26, 27, 28] Embeds traffic rules in state representations and reward functions. Employs hybrid strategies (e.g., imitation learning, formal logic) to improve safety and exploration. Advantages: Adaptive policy learning from dynamic environments. Limitations: Reward engineering complexity, extensive simulation, and sample inefficiency.
LLM & RAG Methods [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] Uses unstructured text and regulatory documents to adapt driving behavior. Translates natural language rules into formal (temporal) logic and generates context-aware decisions. Often combined with multimodal inputs (e.g., DriveGPT4). Advantages: Dynamic adaptation and enhanced interpretability. Limitations: Latency concerns and risk of LLM hallucinations.
Formal Logic-Based Methods [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] Formalizes traffic regulations via temporal logic, ASP, etc. Integrated into verification, MPC, and optimization frameworks for provable safety. Advantages: Strong safety guarantees and high interpretability. Limitations: Worst-case planning time grows exponentially with rule size and challenges with real-time integration.
Hybrid Symbolic & Neural Methods [42] (also in Formal Logic), [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65] [37, 40] (also in LLM) Fuses deep learning–based perception with explicit symbolic reasoning. Incorporates symbolic constraints in loss functions or via LLM–generated rules. Advantages: Balances flexibility with rule compliance. Limitations: Increased integration complexity and synchronization challenges.
Diffusion & Other Learning-Based Approaches [65, 66, 67] Leverages diffusion processes to generate multimodal, rule–compliant trajectories. Blends stochastic exploration with explicit constraint encoding. Advantages: Captures diverse trajectory distributions. Limitations: Balancing exploration and speed with strict rule adherence can be challenging.
@@ II Literature Survey
This section presents a high-level overview of the diverse methodologies for integrating domain knowledge, traffic rules, and commonsense reasoning into AD systems. Our survey is organized into several subsections—ranging from knowledge graph and ontology-based methods, reinforcement learning (RL) methods, and LLM approaches to formal logic-based and hybrid neuro-symbolic methods—that collectively cover the spectrum of current research. To navigate the diverse methodologies, Table I offers a concise, comparative summary of the key studies covered in this survey. The table categorizes each representative study by its methodological approach, emphasizes the main contributions, and outlines strengths and weaknesses. Some works fall under multiple categories; we assign each a primary class but list cross-references. This high-level overview facilitates understanding how these diverse approaches interconnect and serves as a roadmap for the detailed discussions in subsequent sections.
@@ II-A Knowledge-Graph and Ontology-Based Methods
Recent advancements in AD have highlighted the importance of structured knowledge representation using ontologies and knowledge graphs (KGs) to model complex urban environments, enhance prediction accuracy, and enable rule-compliant planning. These frameworks provide explicit semantic representations of static road elements (e.g., lanes, traffic signs), dynamic interactions (e.g., vehicle-pedestrian relationships), and regulatory constraints, enabling interpretable reasoning across heterogeneous data sources.
@@
Early works demonstrated the potential of ontology-based reasoning for driving scenarios. [7] pioneered a scenario-adaptive approach using Web Ontology Language (OWL) to formalize spatial states and vehicle-road relationships, integrating these with hidden Markov models for maneuver prediction. At the same time, their method improves prediction horizons through domain knowledge, and manual ontology engineering limits scalability. Similarly, [8] developed intersection-specific ontologies with rule-based decision-making, and [9] established a framework for querying static and dynamic traffic semantics in real time. These studies underscore the trade-off between interpretability and adaptability: rule-based ontologies excel in structured environments like intersections [8] but struggle with unseen scenarios.
@@
SemanticFormer [10] improves trajectory prediction by considering traffic participants, road topology, and signs, enhancing performance by 4-5% when added to existing trajectory predictors. Hybrid methods combining ontologies with probabilistic models address uncertainty inherent in traffic interactions. [11] merged high-level traffic coordination ontologies with low-level trajectory models, while [12] integrated description logic with probabilistic semantics to handle incomplete information. [13] extended this by encoding interaction probabilities for long-term behavior prediction, though computational delays hindered real-time control. [14] augmented scene understanding by integrating KGs extracted from textual corpora, enabling causal reasoning for novel scenarios. However, offline knowledge curation in their approach limits responsiveness to dynamic environments.
@@
Ontologies also support real-time decision-making through hierarchical reasoning. [15] leveraged driving experience knowledge bases for situation estimation, while [16] modeled critical factors via joint description logic and rule-based reasoners. These methods enhance transparency in regulatory compliance but face scalability challenges as ontology complexity grows [9]. A comprehensive survey by [17] emphasizes that while ontological frameworks improve interoperability and rule awareness, their rigidity complicates adaptation to unexpected events.
@@
Current research focuses on mitigating manual engineering burdens through standardization and automation. Domain-specific ontologies like ASAM OpenXOntology111www.asam.net/standards/asam-openxontology provide reusable templates for traffic concepts, reducing development efforts [18, 19]. Promisingly, LLMs are being explored for semi-automated KG generation [20], potentially enabling dynamic knowledge integration with sensor data. Such advances aim to balance the interpretability of symbolic reasoning with the adaptability of data-driven methods, addressing long-standing challenges in real-time performance and scalability [17, 13].
@@ II-B Reinforcement Learning (RL) Methods
Reinforcement learning (RL) has emerged as a powerful paradigm for developing AD systems, enabling sequential decision-making through dynamic environment interactions. By integrating safety constraints, regulatory compliance, and structured domain knowledge into reward functions and state representations, RL frameworks excel in learning adaptive policies for complex driving scenarios. Recent advancements in this domain focus on curriculum learning (progressively increasing the complexity of training scenarios), safety-centric reward design, and sim-to-real transfer while addressing challenges such as sample inefficiency and reward engineering.
A key trend in RL-based methods is using structured state encodings to embed traffic rules and safety priors. For instance, [21] proposes a graph-based RL framework that employs curriculum learning to progressively increase the complexity of highway driving scenarios. By integrating explicit traffic-rule information into the state representation using graph neural networks, their method accelerates learning and improves collision avoidance and lane-change efficiency in dense traffic. Similarly, [22] addresses the challenge of adapting to evolving traffic regulations by designing a law-adaptive backup policy. Their approach translates natural language traffic laws into formal constraints, enabling the RL agent to recognize potential violations and trigger re-planning. While these methods enhance compliance and safety, they introduce the complexity of reward engineering and policy evaluation.
@@
Bridging the gap between simulation and real-world deployment remains a critical challenge for RL-based planners. [26] investigates the sim-to-real transfer of RL policies, training agents in virtual environments and adapting them to real-world conditions. Their work highlights the potential of RL for robust performance in heterogeneous scenarios but also underscores the challenges of achieving reliable generalization. Surveys such as [27] provide a broader perspective, categorizing RL-based approaches by state representation (e.g., raw sensor inputs vs. graph abstractions) and reward formulations. These surveys emphasize the trade-offs between sample efficiency and safety guarantees, offering insights into the design of scalable and reliable RL frameworks.
Despite their advantages, RL-based methods face several challenges. Reward engineering remains a non-trivial task, as balancing safety, compliance, and efficiency in reward design is critical for policy performance. Sample inefficiency is another significant bottleneck, as complex environments demand extensive training data, prolonging development cycles. The sim-to-real gap further complicates deployment, with discrepancies between simulated and real-world dynamics hindering reliable policy transfer. Additionally, the “black-box” nature of RL policies complicates verification in safety-critical contexts, necessitating interpretable and verifiable representations.
Emerging directions in the RL-based model focus on addressing these challenges. Evolutionary strategies, such as those proposed by [23], offer efficient exploration mechanisms for safe policy learning. Formal methods like linear temporal logic (LTL) [24] provide verifiable compliance guarantees, while hybrid architectures blending imitation learning with RL [25] enhance human-like behavior and adaptability. Recent work in Multi-Agent RL (MARL) using self-play shows promise for achieving robust performance and handling complex interactions in simulated real-world benchmarks [28], potentially addressing some scalability and interaction modeling challenges. Future research should prioritize scalable reward automation, robust sim-to-real pipelines, and interpretable policy representations to advance the real-world applicability of RL-based autonomous driving systems.